Write the equation of the line passing through (4, 4) and (4, 2).
Answer by checkley71(8403)

FIRST YOU NEED TO FIND THE SLOPE (m)=(Y2-Y1)/(X2-X1)
(2-4)/(4-4)=-2/0 = INFINITY.
WHICH IDENTIFIES A VERTICLE LINE THROUGH X=4.
THUS YOU HAVE THE LINE EQUATION X=4.